# Delay-Aware Task Offloading for Heterogeneous VLC-RF-based Vehicular Fog Computing

### v1

@@ I Introduction
For task offloading, bipartite task partitioning divides computing tasks into two components, which are then processed by onboard central processing unit (CPU), edge servers, or dedicated service vehicles [7, 17]. Nevertheless, such a scheme only leveraged the computing resources from a limited number of vehicles, failing to exploit substantial idle resources in the network. In the case of multi-partition offloading, tasks were pre-split and offloaded to multiple idle vehicles based on fuzzy logic [2] and overlapping coalition game [16], yet static task partitioning ignored dynamic variations in achievable data rates and available vehicle computing resources, leaving significant optimization potential. Furthermore, dynamic multi-partitioning of tasks and computing resources trading for task execution were enabled via multi-agent reinforcement learning [14], but its simplified assumption of interference-free parallel transmissions undermines its applicability in realistic multi-vehicle scenarios.
For task transmission, existing VFC implementations predominantly depend on radio frequency (RF) communications for broad coverage, suffering from spectral congestion and competition-induced failures, especially in dense vehicular networks [9]. In contrast, via directional line-of-sight (LoS) links between headlamps/taillights and photodiodes, visible light communication (VLC) inherently mitigates inter-vehicle interference, but the coverage area is spatially constrained [5, 4]. Under this circumstance, the heterogeneous VLC-RF architecture aims to combine the interference resilience of VLC and the extended coverage of RF [10], which is expected to reduce the transmission delay and ensure the timely completion of computing tasks.
In this context, the primary contributions of this paper are summarized as follows. Firstly, to the best of our knowledge, it is the first time to integrate heterogeneous VLC-RF communications into VFC, with a comprehensive system model being constructed, where the dynamic multi-partitioning of tasks is enabled to fully utilize the idle computing resources. Secondly, the average TPD minimization problem is formulated for the proposed heterogeneous VFC system, and the corresponding residual-based majorization-minimization (RBMM) algorithm is designed. Thirdly, numerical results demonstrate a 15% average TPD reduction of the heterogeneous VFC system with the proposed RBMM algorithm compared to VFC systems relying solely on VLC or RF.
@@ II System Model
The task partition matrix is defined as $\bm{M}\in\mathbb{R}^{(V+1)\times U}$, where $\mathbb{R}$ indicates the field of real numbers. Its element $m_{v,u}$ denotes the task proportion offloaded from TV $u$ to SV $v$, and $m_{V+1,u}$ corresponds to the locally processed proportion at TV $u$. Besides, the computing resources allocation matrix is defined as $\bm{F}\in\mathbb{R}^{V\times U}$, with its element $f_{v,u}$ specifying the proportion of computing resources allocated by SV $v$ to process the task of TV $u$. Therefore, the computing delay for SV $v$ to process the task of TV $u$ is calculated by $Z_{v,u}=m_{v,u}\rho_{u}\kappa_{u}/(f_{v,u}C_{v})$. In addition, the LC delay of TV $u$ is given by $Z_{V+1,u}=m_{V+1,u}\rho_{u}\kappa_{u}/C_{u}$.
@@ III Average Delay Minimization Problem
Time slot constraint: Constraint in (7e) guarantees the tasks of all TVs are processed within a time slot.
@@ IV Task Offloading and Resource Allocation
In this section, the non-convex constraint in problem (P1) is transformed to a second-order cone (SOC) constraint, and then the reformulated problem is solved by the proposed RBMM algorithm.
@@ V Simulation Results
Figure 3: Comparison of multiple task offloading methods under various computing resources.
The convergence characteristics of the proposed RBMM algorithm are illustrated in Fig. 2, where progressive reductions in the average TPD are observed through successive iterations, and the average TPD basically converges at 5 iterations. Besides, compared with the initial iteration point, the proposed RBMM algorithm achieves over 30% reduction in the average TPD, with even more reductions observed in maximum TPD, thereby validating the effectiveness of the RBMM algorithm. Notably, both average LC and offloading delays asymptotically converge toward equilibrium, which arises because the TPD of a TV is governed by the maximum value among its LC delay and parallel offloading delays to multiple SVs. On this occasion, the proposed RBMM algorithm effectively reduces TPD by dynamically balancing these delay components, thereby driving their coordinated convergence.
@@
In Fig. 3, the proposed RBMM algorithm is compared with various baselines across varying computing resources. 1) VLC only: Tasks are offloaded only via VLC. 2) RF only: Tasks are offloaded only via RF. 3) LC [11]: All tasks are computed by the corresponding TV locally without offloading. 4) Particle swarm optimization (PSO) [14] for heterogeneous VFC: The high-quality optimization solution is found through the cooperation and information sharing between individuals in the particle swarm. 5) Covariance matrix adaptation-evolution strategy (CMA-ES) [3] for heterogeneous VFC: By dynamically adjusting the covariance matrix to guide the search direction and step size of the evolution strategy, a high-quality optimization solution is achieved.
@@
In Fig. 4, the dual impact of task computation intensity and subchannel bandwidth is elucidated. As demonstrated in Fig. 4(a), when subchannel bandwidth increases, the achievable data rates of both VLC and RF subsystems are improved, thereby reducing the transmission delay and inducing a consistent decline in the average TPD. Conversely, higher task computation intensities necessitate more resources for computing, leading to prolonged computing delay and consequently elevating the average TPD. Notably, under high task computation intensity conditions, subchannel bandwidth expansion yields more pronounced reductions in the average TPD. As the computation intensity increases, computing delay is amplified, which necessitates a larger proportion of tasks offloading from TVs to SVs to balance LC and offloading delays. Under these circumstances, the bandwidth expansion effectively mitigates the transmission delay induced by the increased offloading tasks under high computation intensity, thereby further reducing average TPD. Concurrently, these analytical insights are further validated in Fig. 4(b), where the task partition proportions of some TVs are demonstrated. When the task computation intensity increases from 50 to 250 cycles/bit, a reduction of at least 13% in the proportion of locally computed tasks per TV is observed, where the reduced part of the tasks is offloaded to other SVs, thus effectively reducing the average TPD.
@@ VI Conclusion
In this paper, a heterogeneous VLC-RF-based VFC system was proposed, enabling the dynamic task multi-partitioning and offloading to idle SVs via both VLC and RF. Based on the constructed comprehensive system model, the average TPD minimization problem was formulated and solved by the proposed RBMM algorithm, which jointly optimized task offloading and computing resource allocation. Furthermore, numerical simulations demonstrated that the heterogeneous VFC system with the proposed RBMM algorithm achieved a 15% average TPD reduction compared to VLC-only or RF-only VFC systems. Moreover, the proposed algorithm demonstrated a significant reduction in average TPD relative to baseline methods, highlighting its potential to support delay-sensitive services in next-generation transportation networks.

### v2

@@ I Introduction
For task offloading, bipartite task partitioning divides computing tasks into two components, which are then processed by onboard central processing unit (CPU), edge servers, or dedicated service vehicles [7]. Nevertheless, such a scheme only leveraged the computing resources from a limited number of vehicles, failing to exploit substantial idle resources in the network. In the case of multi-partition offloading, tasks were pre-split and offloaded to multiple idle vehicles based on fuzzy logic [2] and overlapping coalition game [17], yet static task partitioning ignored dynamic variations in achievable data rates and available vehicle computing resources, leaving significant optimization potential. Furthermore, dynamic multi-partitioning of tasks and computing resources trading for task execution were enabled via multi-agent reinforcement learning [14], but its simplified assumption of interference-free parallel transmissions undermines its applicability in realistic multi-vehicle scenarios.
For task transmission, existing VFC implementations predominantly depend on radio frequency (RF) communications for broad coverage, suffering from spectral congestion and competition-induced failures, especially in dense vehicular networks [9]. In contrast, via directional line-of-sight (LoS) links between headlamps/taillights and photodiodes, visible light communication (VLC) inherently mitigates inter-vehicle interference, which was utilized for task offloading in edge computing [16]. Under this circumstance, the heterogeneous VLC-RF architecture aims to combine the interference resilience of VLC and the extended coverage of RF [10], which is expected to reduce the transmission delay and ensure the timely completion of computing tasks.
The primary contributions of this paper are summarized as follows. Firstly, to the best of our knowledge, it is the first time to integrate heterogeneous VLC-RF communications into VFC, with a comprehensive system model being constructed, where the dynamic multi-partitioning of tasks is enabled to fully utilize the idle computing resources. Secondly, the average TPD minimization problem is formulated, and the corresponding residual-based majorization-minimization (RBMM) algorithm is designed. Thirdly, numerical results demonstrate a 15% average TPD reduction of the heterogeneous VFC system with the proposed algorithm compared to VFC systems relying solely on VLC or RF.
@@ II System Model
The task partition matrix is defined as $\bm{M}\in\mathbb{R}^{(V+1)\times U}$, where $\mathbb{R}$ indicates the field of real numbers. Its element $m_{v,u}$ denotes the task proportion offloaded from TV $u$ to SV $v$, and $m_{V+1,u}$ corresponds to the locally processed proportion at TV $u$. Besides, the computing resources allocation matrix is defined as $\bm{F}\in\mathbb{R}^{V\times U}$, with its element $f_{v,u}$ specifying the proportion of computing resources allocated by SV $v$ to process the task of TV $u$. Therefore, the computing delay for SV $v$ to process the task of TV $u$ is calculated by $Z_{v,u}=m_{v,u}\rho_{u}\kappa_{u}/(f_{v,u}C_{v})$.
@@ III Average Delay Minimization Problem
Time slot constraint: Constraint in (7e) guarantees the maximum TPD among TVs do not exceed the duration of one time slot.
@@ IV Task Offloading and Resource Allocation
In this section, the constraint in problem (P1) is transformed to a second-order cone (SOC) constraint, and then the reformulated problem is solved by the proposed RBMM algorithm.
@@ V Simulation Results
Figure 3: Comparison of multiple baseline methods under various computing resources.
The convergence characteristics of the proposed RBMM algorithm are illustrated in Fig. 2, where progressive reductions in the average TPD are observed through successive iterations, and the average TPD basically converges at 5 iterations. Besides, compared with the initial iteration point, the proposed RBMM algorithm achieves over 30% reduction in the average TPD, with even more reductions observed in maximum TPD. Notably, both average LC and offloading delays asymptotically converge toward equilibrium, which arises because the TPD of a TV is governed by the maximum value among its LC delay and parallel offloading delays to multiple SVs. On this occasion, the proposed RBMM algorithm effectively reduces TPD by dynamically balancing these delay components, thereby driving their coordinated convergence.
@@
In Fig. 3, the proposed RBMM algorithm is compared with various baselines across varying computing resources. 1) VLC only: Tasks are offloaded only via VLC. 2) RF only: Tasks are offloaded only via RF. 3) LC [11]: All tasks are computed by the corresponding TV locally without offloading. 4) Particle swarm optimization (PSO) [14] for heterogeneous VFC: The high-quality optimization solution is found through the cooperation and information sharing in the particle swarm. 5) Covariance matrix adaptation-evolution strategy (CMA-ES) [3] for heterogeneous VFC: By dynamically adjusting the covariance matrix to guide the search direction and step size of the evolution strategy, a high-quality solution is achieved.
@@
In Fig. 4, the dual impact of task computation intensity and subchannel bandwidth is elucidated. As demonstrated in Fig. 4(a), when subchannel bandwidth increases, the achievable data rates of both VLC and RF are improved, thereby reducing the transmission delay and inducing a consistent decline in the average TPD. Conversely, higher task computation intensities necessitate more resources for computing, leading to prolonged computing delay and consequently elevating the average TPD. Notably, under high task computation intensity conditions, subchannel bandwidth expansion yields more pronounced reductions in the average TPD. As the computation intensity increases, computing delay is amplified, which necessitates a larger proportion of tasks offloading from TVs to SVs to balance LC and offloading delays. In this case, the bandwidth expansion effectively mitigates the transmission delay induced by the increased offloading tasks under high computation intensity, thereby further reducing average TPD. Concurrently, these analytical insights are further validated in Fig. 4(b), where the task partition proportions of some TVs are demonstrated. When the task computation intensity increases from 50 to 250 cycles/bit, a reduction of at least 13% in the proportion of locally computed tasks per TV is observed, where the reduced part of the tasks is offloaded to other SVs, thus effectively reducing the average TPD. Besides, because offloading delay includes transmission delay in addition to computing delay, the aforementioned balancing mechanism between LC and offloading delays leads to a high proportion of tasks being processed locally.
@@ VI Conclusion
In this paper, a heterogeneous VLC-RF-based VFC system was proposed, enabling the dynamic task multi-partitioning and offloading to idle SVs via both VLC and RF. Based on the constructed comprehensive system model, the average TPD minimization problem was formulated and solved by the proposed RBMM algorithm, which jointly optimized task offloading and computing resource allocation. Furthermore, numerical simulations demonstrated that the heterogeneous VFC system with the proposed RBMM algorithm achieved a 15% average TPD reduction compared to VLC-only or RF-only VFC systems. Moreover, the proposed algorithm demonstrated a significant reduction in average TPD relative to baseline methods, highlighting its potential to support delay-sensitive services in next-generation transportation networks. In future work, allocation of both communication and computing resource can be optimized with task offloading to enable efficient VFC.